how much pure acid should be mixed with 5 gallon of a 20% acid solution in orde to get a 50% acid solution
Found 2 solutions by mananth, jorel1380:
Answer by mananth(16946)

------ percent ---------------- quantity
dilute acid 20----------------5 gallons
pure acid 100----------------x gallons
Total 50 ----------------5+x gallons

20*5+100*x=50(5+x)
100+100 x=250+50x
100x-50x=250-100
50x=150
/50
x=3 gallons 100 % pure acid

Answer by jorel1380(3719)

.50x=1+(x-5)
.50x=x-4
4=x-.5x
.5x=4
x=4/.5=8
8-5=3 gals. of pure acid to be mixed with 5 gals. of 20% acid
5x.20+3=4
8x.5=4